# Towards the Worst-case Robustness of Large Language Models

### v1

@@ 5.1 Formulation of Certified Robustness for LLMs
Inspired by Cohen et al. (2019), we define certified robustness using randomized smoothing as the following functional optimization problem:
@@
Way III: Certifying the output of a language model. Given a language model $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$ and a judgment oracle $\mathcal{O}:\mathcal{V}^{N}\to\{0,1\}$, we construct a smoothed function $g(\bm{x})=\mathbb{E}[\mathcal{O}(f(\bm{z}))]$, which is exactly the probability that $f(\bm{z})$ produces a benign output. If we can show that $g(\bm{x}_{adv})$ is larger than $\tau$, this demonstrates that the output of the language model is safe with at least $\tau$ probability. However, this approach requires a highly accurate or strictly conservative oracle detector $\mathcal{O}$, which is not feasible with most practical detectors.
Since ways II and III are not strictly rigorous, in the main paper, we focus exclusively on certifying a safety detector.
5.2 Certified Robustness on General Distribution
Previous researchers have addressed certified robustness for simple distributions, such as Gaussian distributions (Cohen et al., 2019), masking distributions (with a fixed masking ratio) (Zeng et al., 2023), and synonym distributions (Ye et al., 2020). However, these methods are not applicable to our uniform and absorbing kernels. To address this, we propose a solution for solving the contained optimization problem in Eq. 1 for any smoothing distribution. We begin by reviewing the Fractional Knapsack Problem:
@@ 5.2 Certified Robustness on General Distribution
(The Fractional Knapsack Problem). Given some items, for each item $\bm{z}$, it has a weight $p(\bm{z}|\bm{x})$ and value $-p(\bm{z}|\bm{x}_{adv})$, the goal is to select items (or fractions of items) such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ does not exceed the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in[0,1]$ denotes the fraction of each item chosen.
The solution to the Fractional Knapsack Problem relies on a greedy algorithm: we prioritize items by their value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, selecting items in descending order of this ratio until the capacity $p_{A}$ is reached. This approach is optimal because it maximizes the contribution of each item (or fraction of an item) per unit weight added to the knapsack (Aho & Hopcroft, 1974; Cormen et al., 2022).
Therefore, to solve Eq. 1, we simply enumerate all possible $\bm{z}$, sort them by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ in descending order, and select items until the cumulative weight reaches $p_{A}$, as shown in Algorithm 2. Each time we select a $\bm{z}$, we consume $p(\bm{z}|\bm{x})$ from $p_{A}$, but add $p(\bm{z}|\bm{x}_{adv})$ to $p_{adv}$. Consequently, we refer to the negative value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ as the trading rate. The larger the trading rate, the greater the increase in $p_{adv}$, the better “our trade” is.
Algorithm 2 Fractal Knapsack Solver for Eq. 1
0:  Smoothing distributions $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$, threshold $\tau$, $p_{A}=g(\bm{x})$. Output: Whether $g$ is provably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$.
1:  Compute $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ for each $\bm{z}\in\mathcal{X}$, sort items in descending order.
2:  Initialize weights and value $W,V\leftarrow 0$.
3:  for each item $\bm{z}$ in sorted order do
4:     if $W+p(\bm{z}|\bm{x})\leq p_{A}$ then
5:        $W\leftarrow W+p(\bm{z}|\bm{x})$ and $V\leftarrow V+p(\bm{z}|\bm{x}_{adv})$
6:     else
7:        Select fraction of $\bm{z}$ to fill remaining $p_{A}-W$
8:        $V\leftarrow V+\left(p(\bm{z}|\bm{x}_{adv})\cdot\frac{p_{A}-W}{p(\bm{z}|\bm{x})}\right)$
9:        Break
10:     end if
11:  end for
12:  Return: $\mathbb{I}\{V\geq\tau\}$
@@
(Proof in Sec. C.1). Algorithm 2 exactly solves Eq. 1.
@@
Note that the tightness of this bound relies on the assumption that the hypothesis set of $f$ includes all functions mapping $\mathcal{V}^{N}$ to $[0,1]$. If we restrict the hypothesis set to functions that map to $\{0,1\}$ (i.e., hard functions that output only 0 or 1), this reduces to a 0-1 Knapsack problem. In such cases, a tighter bound can be obtained using dynamic programming. See Sec. C.8 for details.
To apply these solvers to specific smoothing kernels, a brute-force way is to enumerate all possible $\bm{z}$ and perform Algorithm 2 for these $\bm{z}$. This requires $O(|\mathcal{V}|^{N})$ for Fractal Knapsack and $O(|\mathcal{V}|^{2N})$ for 0-1 Knapsack. A clever way to do this is to categorize these items according to value-to-weight ratios and calculate the number of items (volume) for each category. Formally, the volume $v(\gamma)$ for trading rate $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma$ is defined as:
$v(\gamma)=\sum_{\bm{z}}p(\bm{z}|\bm{x})\mathbb{I}\left\{\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma\right\}.$
In Sec. 5.3 and Sec. 5.4, we show how to reduce the time complexity from $O(|\mathcal{V}|^{N})$ to $O(1)$ for absorbing kernel and from $O(|\mathcal{V}|^{N})$ to $O(2d)$ for uniform kernel.
@@ 5.3 Certified Robustness on Absorbing Kernel
For simplicity, we use the subscript $p$ to denote the common part of $\bm{x}$ and $\bm{x}_{adv}$, and the subscript $s$ to denote the differing part of $\bm{x}$ and $\bm{x}_{adv}$. Clearly, $\bm{x}_{p}=\bm{x}_{adv_{p}}$. We denote $\beta$ as the probability of turning each word into a masked token, and $\bar{\beta}=1-\beta$ as the probability of remaining unchanged.
To solve Eq. 1 for absorbing kernel, it suffices to calculate all trading rates $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ and their corresponding volumes. We give these results in the following theorem:
@@
(Certified robustness for Absorbing Kernel, proof in Sec. C.3) Divide $\mathcal{V}^{N}$ into $L_{1}$ and $L_{2}$ that $L_{1}\cup L_{2}=\mathcal{V}^{N}$ and $L_{1}\cap L_{2}=\emptyset$, where $L_{1}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are all masked tokens}\}$, $L_{2}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are not all masked tokens}\}$. Clearly, we have the trading rate:
@@
By plugging these results into Algorithm 2, we can compute the certified bound using the absorbing kernel. Interestingly, due to the simplicity of the trading rate and volume, we can derive an analytic solution for the certified radius as $\lfloor\frac{\ln(1-p_{A}+\tau)}{\ln\beta}\rfloor$. For details, refer to Sec. C.4.
Table 2: The average certified $\ell_{0}$ radius (at the false positive rate 0) against $\ell_{0}$ attack.
Vicuna-7B Llama2-7B Llama3-8B GPT-4o Trivial Bound Bayesian Bound
DiffTextPure (Absorb) 1.00 1.92 1.82 2.00 2.12 2.10
DiffTextPure (Uniform) 1.02 1.86 1.54 2.02 2.12 2.10
@@ 5.4 Certified Robustness on Uniform Kernel
For value-to-weight ratio:
For its volume:
With a slight abuse of notation, we still use $\bar{\beta}$ as the probability of each token remaining unchanged, $\beta=1-\bar{\beta}$ as the probability of doing perturbation, $\alpha=\frac{\beta}{|\mathcal{V}|-1}$ as the probability of perturbing into each of the other tokens.
@@
Proof.
Here, $i$ represents the number of perturbed tokens in the normal suffix, and $j$ represents those in the adversarial suffix. Only when $i+j\geq d$ can $\bm{z}$ be derived from both $\bm{x}$ and $\bm{x}_{adv}$. Among the $i+j$ modified tokens, $i+j-d$ tokens must be perturbed to arbitrary tokens, excluding the counterparts in either the normal or adversarial suffixes, resulting in $(|\mathcal{V}|-2)^{i+j-d}$ possible values for these tokens. The $i$ tokens can be chosen from any positions in the suffix, giving the term $\binom{d}{i}$. Among these $i$ tokens, $d-j$ must match the adversarial suffix, while the remaining $i+j-d$ can be perturbed to any of $|\mathcal{V}|-2$ possible values, leading to a total arrangement count of $\binom{i}{d-j}$. ∎
A notable property of the uniform kernel is that if $g(\bm{x})=1$, the certified radius becomes infinite. This occurs because the support of $p(\bm{z}|\bm{x})$ spans the entire space $\mathcal{V}^{N}$. When $g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=1$, it implies that $f(\bm{z})=1$ for all $\bm{z}$. Consequently, $g(\bm{x}_{adv})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ will also equal 1, which is always greater than $\tau$. In contrast, with the absorbing kernel, $g(\bm{x}_{adv})$ cannot exceed $\beta^{d}$. Therefore, the certified radius is at most $\frac{\log\tau}{\log\beta}$. From this perspective, the uniform kernel closely resembles the behavior of the Gaussian distribution in the image domain, where the certified radius can also potentially be infinite.
In the following theorem, we theoretically demonstrate that the uniform kernel outperforms the absorbing kernel not only when $p_{A}=1$ but across all values of $p_{A}$.
@@
(Proof in Sec. C.5.) The certified radius of uniform kernel is always larger than absorbing kernel given the same accuracy $p_{A}$, threshold $\tau$, perturbing probability $\beta$, i.e.,
$\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V})\geq\text{certify}(\text{absorb},p_{A},\tau,\beta).$
The equality holds when $|\mathcal{V}|\to\infty$.
5.5 Upper Bound of Certified Radius
Definition 5.8.
where $\mathcal{O}$ represents the judgment oracle.
Trivial Bound. Changing $K(\bm{x})$ words will alter the semantics of the input. Therefore, we can certify at most $\ell_{0}$ attacks involving $K(\bm{x})-1$ words, i.e.,
$R(\bm{x})\leq K(\bm{x})-1.$

### v2

@@ 4.1 Formulation of Certified Robustness for LLMs
Inspired by Cohen et al. [23], we define certified robustness using randomized smoothing as the following functional optimization problem:
@@
Way III: Certifying the Output of a Language Model. Given a language model $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$ and a judgment oracle $\mathcal{O}:\mathcal{V}^{N}\to\{0,1\}$, we construct a smoothed function $g(\bm{x})=\mathbb{E}[\mathcal{O}(f(\bm{z}))]$ (i.e., returning 1 when the output is safe and 0 when unsafe), which represents the probability that $f(\bm{z})$ produces a benign output. If we can show that $g(\bm{x}_{adv})$ is greater than $\tau$, this demonstrates that the output of the language model is safe with at least probability $\tau$. This definition is general, as the judgment oracle can encompass other benchmarks, enabling certification of various desired properties (e.g., coding, math, CoT, safety). However, although we obtain a tight lower bound for Eq. 1, we may still be unable to derive a practical bound for this definition. This limitation may be addressed in the future by incorporating additional neural network-dependent constraints. See Sec. I.1 for details.
Therefore, in the main paper, we focus exclusively on certifying a safety detector.
4.2 Certified Robustness on Bounded $f$
Previous researchers have addressed certified robustness for simple distributions, such as Gaussian distributions [23], masking distributions (with a fixed masking ratio) [29], and synonym distributions [28]. However, these methods are not applicable to a general distribution. To address this, we propose a solution for solving the constrained optimization problem in Eq. 1 for any smoothing distribution. We begin by reviewing the Fractional Knapsack Problem:
@@ 4.2 Certified Robustness on Bounded $f$
(The Revised Fractional Knapsack Problem). Given a set of items, each item $\bm{z}$ has a weight $p(\bm{z}|\bm{x})$ and a value $-p(\bm{z}|\bm{x}_{adv})$. The goal is to select items (or fractions of items) such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in[0,1]$ denotes the fraction of each item chosen.
There are two differences between Definition 4.2 and the traditional fractional knapsack problem. First, Definition 4.2 is a minimization problem rather than a maximization problem, but this can be addressed by defining the value of each item as $-p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, Definition 4.2 requires that the total weight of items must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it. Since the greedy algorithm of fractional knapsack solvers always finds a solution that precisely fits the knapsack (as demonstrated in Algorithm 1), this constraint does not pose an issue.
The solution to the Fractional Knapsack Problem relies on a well-known greedy algorithm: we prioritize items by their value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, selecting items in descending order of this ratio until the capacity $p_{A}$ is reached. This approach is optimal because it maximizes the contribution of each item (or fraction of an item) per unit weight added to the knapsack [50, 51].
Therefore, to solve Definition 4.2, we simply enumerate all possible $\bm{z}$, sort them by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ in descending order, and select items until the cumulative weight reaches $p_{A}$, as shown in Algorithm 1. Each time we select a $\bm{z}$, we consume $p(\bm{z}|\bm{x})$ from $p_{A}$, but add $p(\bm{z}|\bm{x}_{adv})$ to $p_{adv}$. Consequently, we refer to the negative value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ as the trading rate. The larger the trading rate, the greater the increase in $p_{adv}$, the better “our trade” is.
Algorithm 1 Fractal Knapsack Solver for (1)
0:  Smoothing distributions $p(\bm{z}|\bm{x})$, $p(\bm{z}|\bm{x}_{adv})$, threshold $\tau$, $p_{A}=g(\bm{x})$.
0:  $g$ is robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$.
1:  Sort $\bm{z}\in\mathcal{X}$ by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ (descending), and initialize $W,V\leftarrow 0$.
2:  for each $\bm{z}$ in sorted order do
3:     if $W+p(\bm{z}|\bm{x})\leq p_{A}$ then
4:        $W\leftarrow W+p(\bm{z}|\bm{x})$, $V\leftarrow V+p(\bm{z}|\bm{x}_{adv})$.
5:     else
6:        Select fraction of $\bm{z}$ to fill remaining $p_{A}-W$ by $V\leftarrow V+\left(p(\bm{z}|\bm{x}_{adv})\cdot\frac{p_{A}-W}{p(\bm{z}|\bm{x})}\right)$
7:     end if
8:  end for
@@
(Proof in Sec. C.1 and [50]). Algorithm 1 exactly solves the functional minimization part in Eq. 1.
After solving the functional minimization $\min_{f}$, one can simply enumerate all $\bm{x}_{adv}$ and run the knapsack solver for each $\bm{x}_{adv}$, identifying the minimal one for solving the input minimization problem $\min_{\bm{x}_{adv}}$. However, it is usually not necessary to enumerate all $\bm{x}_{adv}$ due to an equivalence: if this $f$ performs worst on a given $\bm{x}_{adv}$, then there exists another $f$ that performs worst on a different $\bm{x}_{adv}$. In other words, the functional minimization $\min_{f}$ already implicitly solves most of the input minimization $\min_{\bm{x}_{adv}}$.
4.3 Certified Robustness on Binary $f$
Note that the tightness of Algorithm 1 relies on the assumption that the hypothesis set of $f$ includes all functions mapping $\mathcal{V}^{N}$ to $[0,1]$. If we restrict the hypothesis set to functions that map to $\{0,1\}$ (i.e., hard functions that output only 0 or 1), this reduces to a 0-1 Knapsack problem, yielding a tighter result.
Definition 4.4.
(The Revised 0-1 Knapsack Problem). Given a set of items, for each item $\bm{z}$, it has a weight $p(\bm{z}|\bm{x})$ and a value $-p(\bm{z}|\bm{x}_{adv})$. The goal is to select items such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in\{0,1\}$ indicates whether each item is chosen.
There are still two differences between Definition 4.4 and the traditional 0-1 knapsack problem. First, the minimization problem can still be converted to a maximization problem by defining the value of each item as $-p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, the requirement that the total weight must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it, introduces additional complexity. While the traditional 0-1 knapsack problem can be reduced to this problem by introducing a slack variable, this problem cannot be reduced to the traditional 0-1 knapsack problem (as it requires an additional constraint). In other words, this problem is more challenging than the traditional 0-1 knapsack problem. Fortunately, we can still devise a dynamic programming approach to solve it; see Sec. C.2 for details.
@@ 4.3 Certified Robustness on Binary $f$
Note that the certified robustness obtained by these two solvers is indeed tight, i.e., there exists a worst-case bounded function $f$ for the fractional knapsack solver and a worst-case binary function $f$ for the 0-1 knapsack solver that satisfy all constraints in Eq. 1, with $g(\bm{x}_{adv})$ equal to the lower bound obtained by our solvers. In other words, the bound for Eq. 1 cannot be further improved. In the future, one might modify Eq. 1 to introduce additional constraints on the base model $f$ (e.g., Lipschitz continuity [52, 53]) to obtain a tighter bound.
5 Case Studies
In this section, we conduct two case studies, analyzing the certified robustness using two popular smoothing kernel $p(\bm{z}|\bm{x})$ —- a uniform kernel (i.e., the forward distribution in diffusion models [54, 20]) and an absorbing kernel (i.e., the forward distribution in mask generation [55, 56]). We show that when they achieve the same standard accuracy, the robustness of the former is strictly greater than that of the latter (and they are equal when the vocabulary size $|\mathcal{V}|\to\infty$).
@@ 5.1 Certified Robustness on Absorbing Kernel
Definition 5.1.
(Absorbing Kernel). An absorbing kernel perturbs each token independently. Each token is replaced with a special masked token $[M]$ with probability $\beta$, and remains unchanged with probability $\bar{\beta}=1-\beta$:
$p(\bm{z}_{i}|\bm{x}_{i})=\begin{cases}\bm{x}_{i}&\text{w.p.\ }\beta,\\ \text{[M]}&\text{w.p.\ }\bar{\beta}=1-\beta.\end{cases}$ (2)
For simplicity, we use the subscript $p$ to denote the common part of $\bm{x}$ and $\bm{x}_{adv}$, and the subscript $s$ to denote the differing part of $\bm{x}$ and $\bm{x}_{adv}$. Clearly, $\bm{x}_{p}=\bm{x}_{adv_{p}}$.
(a) Absorbing kernel.
(b) Uniform kernel.
(c) Different vocabulary size.
To apply fractional knapsack solvers to specific smoothing kernels, a brute-force approach is to enumerate all possible $\bm{z}$ and perform Algorithm 1 for each $\bm{z}$. However, fractional knapsack solvers only depend on the value-to-weight ratio and the total weight of items with a given value-to-weight ratio. If multiple items share the same value-to-weight ratio, we can group these items into categories and calculate the total weight (volume) for each category. Formally, the volume $v(\gamma)$ for a trading rate $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma$ is defined as:
$v(\gamma)=\sum_{\bm{z}}p(\bm{z}|\bm{x})\mathbb{I}\left\{\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma\right\}.$ (3)
This approach not only significantly reduces the time complexity but also provides a clearer understanding of the relationship between $p_{adv}:=g(\bm{x}_{adv})$ and $p_{A}:=g(\bm{x})$.
We provide these results for the absorbing kernel in the following theorem:
@@
(Proof in Sec. D.2) Divide $\mathcal{V}^{N}$ into $L_{1}$ and $L_{2}$ that $L_{1}\cup L_{2}=\mathcal{V}^{N}$ and $L_{1}\cap L_{2}=\emptyset$, where $L_{1}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are all masked tokens}\}$, $L_{2}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are not all masked tokens}\}$. Clearly, we have the trading rate:
@@
By applying these results to Algorithm 1, we show that for the absorbing kernel, if $p_{A}=g(\bm{x})\leq 1-\beta^{d}$, no robustness guarantee can be obtained. For $p_{A}\geq 1-\beta^{d}$, we can obtain a robustness guarantee that $p_{adv}=g(\bm{x}_{adv})\geq p_{A}-(1-\beta^{d})$, with a maximum of $\beta^{d}$, as illustrated in Figure 1(a).
Table 2: The average certified $\ell_{0}$ radius against $\ell_{0}$ attack.
Vicuna-7B Llama2-7B Llama3-8B GPT-4o Human
Absorb 1.00 1.92 1.82 2.00 2.12
Uniform 1.02 1.86 1.54 2.02 2.12
SmoothLLM 2.25 3.24 3.16 3.84 4.04
@@ 5.2 Certified Robustness on Uniform Kernel
Definition 5.3.
(Uniform Kernel). A uniform kernel perturbs each token independently. Each token is replaced with any other token in the vocabulary $\mathcal{V}$ with probability $\alpha=\frac{\beta}{|\mathcal{V}|-1}$, and remains unchanged with probability $\bar{\beta}=1-\beta$:
$p(\bm{z}_{i}|\bm{x}_{i})=\begin{cases}\bm{x}_{i}&\text{w.p.\ }\bar{\beta}=1-\beta,\\ \bm{v}\in\mathcal{V}\setminus\{\bm{x}_{i}\}&\text{w.p.\ }\alpha=\frac{\beta}{|\mathcal{V}|-1}.\end{cases}$ (4)
We provide the volume for each value-to-weight ratio of uniform kernel in the following theorem:
@@
A notable property of the uniform kernel is that if $g(\bm{x})=1$, then $g(\bm{x}_{adv})$ is also one. This occurs because the support of $p(\bm{z}|\bm{x})$ spans the entire space $\mathcal{V}^{N}$. When $g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=1$, it implies that $f(\bm{z})=1$ for all $\bm{z}$. Consequently, $g(\bm{x}_{adv})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ will also equal 1. In contrast, with the absorbing kernel, $g(\bm{x}_{adv})$ cannot exceed $\beta^{d}$. From this perspective, the uniform kernel closely resembles the behavior of the Gaussian distribution in the image domain, where the certified radius can also potentially be infinite [23, 45].
More interestingly, as $|\mathcal{V}|$ increases, the $p_{adv}-p_{A}$ graph of the uniform kernel shifts downward and to the right, and when $|\mathcal{V}|\to\infty$, the $p_{adv}-p_{A}$ graph of the uniform kernel converges to that of the absorbing kernel, as stated in the following theorem:
@@
(Proof in Sec. D.4.) The certified radius of the uniform kernel is always greater than or equal to that of the absorbing kernel given the same accuracy $p_{A}$, threshold $\tau$, and perturbation probability $\beta$, i.e.,
$\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V})\geq\text{certify}(\text{absorb},p_{A},\tau,\beta).$ (6)
Equality holds when $|\mathcal{V}|\to\infty$.